# 3D-Mol: A Novel Contrastive Learning Framework for Molecular Property Prediction with 3D Information

### v2

@@ 4.3 Result
c) Ablation study. To validate the efficacy of our pretraining task, we study the performance of the 3D-Mol encoder in three scenarios: 3D-Mol, 3D-Mol without pretraining, and 3D-Mol without the dihedral-angle graph. The results are shown in Table 3. Compared with the 3D-Mol and 3D-Mol without pretraining, the former performs better in all datasets, demonstrating that our pretraining method can improve encoder performance. Similarly, Compared with the 3D-Mol and 3D-Mol without $G_{d-a}$, the former also shows better performance in all datasets, indicating that the dihedral-angle graph contributes to improved encoder performance. In general, our pretraining and modeling methods enhance the 3D-Mol encoder performance, as the model can more effectively learn the 3D structural information of molecule.
5 Conclusion
@@
In this paper, we propose 3D-Mol, a novel framework for molecular property prediction. It leverages a unique hierarchical graph-based embedding to fully capture 3D structural information of molecules. This model is further enhanced by a contrastive learning component using 3D conformational data to identify structural similarities. Demonstrating superior performance over seven benchmark models, 3D-Mol shows great potential in enhancing AI-assisted drug discovery and molecular design.
@@ Acknowledgment
Appendix A Self-supervised Tasks Based on Physicochemical and Geometric Properties
$L_{i,j}^{l}=(f_{l}(Fn^{mk}_{n,i},Fn^{mk}_{n,j})-l_{i,j})^{2}$ (16)
where $f_{\phi}(.)$, $f_{\theta}(.)$ and $f_{l}$ are the MLPs for the local geometry task, and $L_{i,j}^{l}$, $L_{i,j,k}^{\theta}$, $L_{i,j,k,p}^{\phi}$ and $L_{i}^{FP}$ are the loss functions for each task. $CE(.)$ is the cross entropy loss, and $bin(·)$ is used to discretize the bond angle and dihedral angle. $Fn^{mk}_{n,i}$ is the latent vector of node i after masking the corresponding sampled items in each task.
$L_{i}^{FP}=BCE(f_{FP}(F^{mk}),FP_{i})$ (19)
$L_{i,j}^{dist}=(f_{dist}(Fn^{mk}_{n,i},Fn^{mk}_{n,j})-dist_{i,j})^{2}$ (20)
where $f_{FP}$ and $f_{dist}(.)$ are the MLPs for global geometric tasks, and $L_{i}^{FP}$ and $L_{i,j}^{dist}$ are the loss functions for each task. $BCE(.)$ is binary cross entropy loss. $F^{mk}$ is the latent vector of the masking molecule.
Appendix B 3D Conformation Descriptor and Fingerprint
B.1 Fingerprint
@@
Appendix C Downstream Dataset
In the finetuning stage, we employed 7 molecular datasets obtained from MoleculeNet. These datasets encompass a range of biophysics datasets such as BACE, physical chemistry datasets like ESOL, and physiology datasets like Tox21. The details are outlined below:
$\bullet$ BACE. The BACE dataset provides both quantitative (IC50) and qualitative (binary label) binding results for a set of inhibitors targeting human $\beta$-secretase 1 (BACE-1).
$\bullet$ Tox21. The Tox21 initiative aims to advance toxicology practices in the 21st century and has created a public database containing qualitative toxicity measurements for 12 biological targets, including nuclear receptors and stress response pathways.
$\bullet$ Toxcast. ToxCast, an initiative related to Tox21, offers a comprehensive collection of toxicology data obtained through in vitro high-throughput screening. It includes information from over 600 experiments and covers a large library of compounds.
$\bullet$ SIDER. The SIDER database is a compilation of marketed drugs and their associated adverse drug reactions (ADRs), categorized into 27 system organ classes.
$\bullet$ ESOL. The ESOL dataset is a smaller collection of water solubility data, specifically providing information on the log solubility in mols per liter for common organic small molecules.
$\bullet$ FreeSolv. The FreeSolv database offers experimental and calculated hydration-free energy values for small molecules dissolved in water.
$\bullet$ Lipo. Lipophilicity is a crucial characteristic of drug molecules that affects their membrane permeability and solubility. The Lipo dataset contains experimental data on the octanol/water distribution coefficient (logD at pH 7.4).

### v3

@@ 4.3.3 Ablation study
To validate the efficacy of our pretraining task, we study the performance of the 3D-Mol encoder in three scenarios: 3D-Mol, 3D-Mol without pretraining, and 3D-Mol without the dihedral-angle graph. The results are shown in Table 3. Compared with the 3D-Mol and 3D-Mol without pretraining, the former performs better in all datasets, demonstrating that our pretraining method can improve encoder performance. Compared to the version without contrastive learning weights from fingerprints and 3D descriptors, 3D-Mol demonstrates superior performance across all datasets. This improvement shows that using weights from fingerprints and 3D descriptors effectively optimizes our contrastive loss, enhancing encoder performance. Similarly, Compared with the 3D-Mol and 3D-Mol without $G_{d-a}$, the former also shows better performance in all datasets, indicating that the dihedral-angle graph contributes to improved encoder performance. In general, our pretraining and modeling methods enhance the 3D-Mol encoder performance, as the model can more effectively learn the 3D structural information of molecule.
Figure 5: Case study of 3D information enhance using 3D-Mol for the BACE task. This figure illustrates the prediction results for three molecules identified as BACE inhibitors, showcasing the efficacy of 3D-Mol, which utilizes three-dimensional molecular information, versus GIN, which relies on two-dimensional data. Each molecule is displayed with its respective name (Q27467123, SCHEMBL12917066, Q27455563), and the results indicate that 3D-Mol accurately predicts BACE inhibition across all cases, while GIN fails to do so.
Figure 6: Case study of 3D information enhance using 3D-Mol for the Freesolv task. This figure illustrates the prediction results for hydration free energy of three molecules, showcasing the efficacy of 3D-Mol, which utilizes three-dimensional molecular information, versus GIN, which relies on two-dimensional data. Each molecule is displayed with its respective name (2,3,4,5-Tetrachlorobiphenyl, p-Anisidine, 3-Nitroaniline), and the results indicate that 3D-Mol accurately predicts hydration free energy across all cases, while GIN fails to do so.
4.3.4 Case Study
In this case study, we explore the predictive capabilities of 3D-Mol, which utilizes three-dimensional molecular data, compared to GIN, which relies solely on two-dimensional information, focusing on their efficacy in identifying inhibitors of the $\beta$-secretase 1 (BACE) enzyme, a crucial target for Alzheimer’s disease treatment. We specifically analyze three molecules: Q27467123, which features a chlorinated aromatic ring that may form key halogen bonds within the BACE active site; SCHEMBL12917066, with a fluorinated aromatic ring that enhances molecular stacking interactions; and Q27455563, a molecule with a bicyclic structure potentially forming multiple hydrogen bonds. As shown in Figure 5, 3D-Mol’s approach, which integrates these 3D conformations and potential interactions such as hydrogen bonding, hydrophobic contacts, and precise geometric fitting into the BACE active site, leads to accurate predictions of their inhibitory activities. In contrast, GIN, unable to account for such intricate 3D-dependent interactions, fails to recognize these molecules as potential inhibitors, demonstrating limitations in capturing the necessary depth and spatial relationships that are crucial for binding. This case study highlights the importance of incorporating 3D spatial awareness in computational models for drug discovery, particularly when targeting enzymes like BACE, where the precise alignment and interaction of molecules within a complex three-dimensional space significantly influence their therapeutic efficacy.
Having validated our 3D-Mol framework with the BACE dataset, we extend its application to the Freesolv dataset, which is instrumental for predicting the hydration free energy of small molecules in water—a crucial determinant for solubility in drug discovery and molecular design. We present a case study of three molecules with distinctive 3D structures that influence their solvation behaviors: 2,3,4,5-Tetrachlorobiphenyl, noted for its hydrophobicity due to chlorine substitution; p-Anisidine, which can engage in hydrogen bonding; and 3-Nitroaniline, where the nitro group impacts solvation through electron withdrawal. Figure 6 displays a comparative analysis of the 3D-Mol and GIN model predictions for these molecules. The Freesolv values and RMSE scores indicate 3D-Mol’s superior ability to capture the impact of molecular geometry on solvation. This case study highlights the necessity of 3D information for precise prediction of solvation-related molecular properties, reinforcing the utility of the 3D-Mol framework in computational chemistry.
5 Conclusion and Discussion
@@
3D-Mol, our novel framework introduced in this paper, significantly advances molecular property prediction by leveraging a unique hierarchical graph-based embedding and a contrastive learning component. This approach allows for a comprehensive capture of 3D molecular structures, setting a new standard in the field. Demonstrating superior performance over multiple benchmark models, 3D-Mol holds immense potential in revolutionizing AI-assisted drug discovery and molecular design.
3D-Mol’s encoder distinctly outperforms all baselines, with performance improvements of 2% in classification and 11% in regression tasks compared to traditional and non-pretraining methods. These advancements are pivotal in drug discovery, as they promise to expedite the development of new treatments through more accurate molecular predictions. The primary challenge faced by 3D-Mol is the time-intensive generation of 3D conformations and the encoding of hierarchical graphs. Future work will focus on optimizing these processes to enhance the model’s efficiency and practical scalability, further solidifying its role in advancing computational chemistry and pharmaceutical development.
@@ Acknowledgment
Appendix A 3D Conformation Descriptor and Fingerprint
A.1 Fingerprint
@@
Appendix B The contribution of pretraining method
Table 4: The contribution of pretraining method. We study the performance of 3D-Mol in three scenarios: contrastive learning only, supervised pretraining only, complete pretraining method, then mark the best results in bold and underline the second best.
Datasets BACE SIDER Tox21 ToxCast ESOL FreeSolv Lipophilicity
# Molecules 1513 1427 7831 8597 1128 643 4200
# Tasks 1 27 12 617 1 1 1
$\rm Contrastive-Learning-Only$ $\rm 0.847_{0.002}$ $\rm\underline{0.652}_{0.012}$ $\rm 0.791_{0.008}$ $\rm 0.693_{0.003}$ $\rm 0.802_{0.036}$ $\rm 1.682_{0.86}$ $\rm 0.616_{0.023}$
$\rm Supervised-Pretraining-Only$ $\rm\underline{0.862}_{0.006}$ $\rm 0.647_{0.007}$ $\rm\textbf{0.796}_{0.002}$ $\rm\underline{0.697}_{0.003}$ $\rm\underline{0.795}_{0.022}$ $\rm\underline{1.664}_{0.070}$ $\rm\underline{0.613}_{0.004}$
$\rm Complete-Pretraining-Method$ $\rm\textbf{0.872}_{0.004}$ $\rm\textbf{0.658}_{0.003}$ $\rm\underline{0.792}_{0.003}$ $\rm\textbf{0.701}_{0.003}$ $\rm\textbf{0.782}_{0.008}$ $\rm\textbf{1.617}_{0.050}$ $\rm\textbf{0.600}_{0.015}$
In this section, we discuss the contributions of contrastive learning and supervised pretraining methods to our pretraining approach. We pretrained our model using three approaches: contrastive Learning only, supervised pretraining only, and complete pretraining method. We compared their performance on 7 benchmark datasets. As the Table 4 shown, the contributions of both contrastive learning and supervised pretraining were less significant than the complete method. These findings emphasize that while both contrastive learning and supervised pretraining contribute positively to the model’s performance, their combination is crucial for achieving optimal results.
Appendix C Finetuning Details
During finetuning for each downstream task, we randomly search the hyper-parameters to find the best performing setting on the validation set and report the results on the test set. Table 5 lists the combinations of different hyper-parameters.
Table 5: hyper-parameter setting
Name Description Range
$lr_{MLP}$ Initial learning rate for MLP head $\{0.004,0.001,0.0004\}$
$lr_{ENC}$ Initial learning rate for the pre-trained encoder $\{0.001,0.0004,0.0001\}$
$Epoch$ The number of epoch in finetuning stage $\{60,80,100\}$
$num_{layer}$ Number of hidden layers in MLP $\{2,3\}$
$Dropout$ Dropout ratio for the model $\{0,0.1,0.2,0.5\}$
$Hidden_{size}$ Size of hidden layers in MLP $\{32,64,128,256\}$